OpenStudy (albert0898):

1. How many grams of solution are needed to obtain 40g K2CO3 from a 10% (aq) solution? 2. How many grams of chlorine are present in a 1L of a 650 PPM (aq) solution?

OpenStudy (anonymous):

mass of 1 mole of K2CO3=2*39+12+3*16=78+12+48=138 mass of 1 mole of water(H2O)=2*1+16=18 i don't remember more than that. i have no reference book with me. i have read about 40 years back.

OpenStudy (aaronq):

no need for moles here. The 10% is probably by mass (though this should really be specified). This means that in 100 g of solution, 10 of those are potassium carbonate. if you need 40 g, then you need 400g of solution. for number two use the fact that 1 ppm = 1 mg/mL, thus the density is 650 mg/mL.